OpenStudy (anonymous):

The volume of a cylinder is 93 cubic feet. What is the volume of a cone that fits exactly inside the cylinder? Provide an explanation and proof for your answer.

OpenStudy (anonymous):

vvhat are you saying nin?

OpenStudy (anonymous):

ah, you should probably explain that vvhhen saying "half" then "1/3"

OpenStudy (e.mccormick):

@lowcard2 Know the two volume formulas involved?

OpenStudy (anonymous):



OpenStudy (solomonzelman):

Since the volume of the cubic cylinder is 93, and knowing that the cone is half of size of the cylinder, therefore the cone is going to be 93 divided by 2, or 46.5.

OpenStudy (solomonzelman):

@lowcard2, do you see my point?

OpenStudy (anonymous):

I don't think you're right @SolomonZelman The volume a right cylinder is given by \[V_{cyl}=\pi R^2 H\] whereas the volume of the cone is \[V_{cone} = \frac{1}{3} \pi R^2 H\] So the volume of the cone would be one third of the volume of the cylinder, \[ V_{cone} = 31 \ ft^3\]

OpenStudy (solomonzelman):

Oh, dang, area of the cirlce, my bad!

OpenStudy (anonymous):

^_^

OpenStudy (solomonzelman):

(•‿•) math is not my thing.